OpenStudy (albert0898):

Mike calculates that from his day's intake of 2,156 calories, four times as many calories were from carbohydrates as from protein, and twice as many calories were from fat as from protein. How many calories were from carbohydrates? **Please Be Descriptive In Your Answer**

OpenStudy (lgbasallote):

do you have an answer?

OpenStudy (anonymous):

4*p=c c+p=2156

OpenStudy (anonymous):

Two equations two unknowns. You have enough information to solve for c and p

OpenStudy (albert0898):

Huh? Please be more thorough. I don't get it.

OpenStudy (albert0898):

Explain all steps please.

OpenStudy (anonymous):

I'm not going to solve for you sorry. I gave you two equations where c=carbohydrates p=protein

OpenStudy (anonymous):

Getting the two equations was the hard part. Everything else should be easy. just solve for c or p. You can do that by substitution method.

OpenStudy (albert0898):

Equation: 2156 = 4c + c + 2c From there: 2156 = 7c And from this you can work out that: c = 308 And because carbohydrates = 4c (c being amount of protein), there must be 1232 calories from carbohydrate. Thanks anyways! xD